# Air-HOLP: Adaptive Regularized Feature Screening for High Dimensional Data

### v1

@@ 5 Application to Prostate Cancer Data
Figure 4: Maximum Multiple R vs. model size using the screened genes by the competing methods Air-HOLP, Ridge-HOLP, and SIS.
@@ 6 Computational Complexity
Air-HOLP shares the same time complexity as Ridge-HOLP. However, Air-HOLP employs an iterative process to update the initial penalty parameter. We investigate how this iterative process impacts the speed of Air-HOLP.
@@
(a) The logarithm of execution time in milliseconds vs. number of features $p$.
(b) Execution time of Air-HOLP vs. Ridge-HOLP in seconds.
Figure 5: Compression of the execution time between Ridge-OLS, Ridge-HOLP, and Air-HOLP. On the left, we present the trimmed mean of the logarithm of execution time in milliseconds for $10$ simulations ($10\%$ trimming). On the right, we present a boxplot of the execution time in seconds for $10$ simulations.
@@ 7 Discussion
Traditional feature screening methods such as SIS measure the marginal relations between features and the response, which can fail when features are correlated. Ridge-HOLP was developed to address this issue and we have now further built on Ridge-HOLP by the development of Air-HOLP, which incorporates adaptively selecting the ridge penalty. This adaptive selection aims to minimize prediction error, thereby effectively balancing the bias-variance trade-off and enhancing feature screening performance.
Through extensive simulation studies, we demonstrated that Air-HOLP consistently outperforms Ridge-HOLP with a fixed penalty parameter across various simulation settings. Air-HOLP achieves higher sure screening probabilities and lower sure screening thresholds. The advantage of Air-HOLP is more pronounced when the number of samples $n$ is closer to the number of features $p$. Additionally, both Air-HOLP and Ridge-HOLP outperformed SIS in correlated settings, making Air-HOLP especially useful for identifying relevant features in high-dimensional correlated data. Moreover, Air-HOLP has the same time complexity as Ridge-HOLP and also enjoys being computationally efficient. We also found that in most simulations, Air-HOLP converged as quickly as in $3$ to $6$ iterations. Therefore, we set the maximum number of iterations to be $10$ in our code.
We further applied Air-HOLP, HOLP and SIS to a prostate cancer genetic dataset. The results confirmed the good performance of Air-HOLP.
@@ Acknowledgments
The authors thank Dr. Helia Farhood for her assistance and support in writing this paper. Samuel Muller was supported by the Australian Research Council Discovery Project Grant (DP230101908).

### v2

@@ 5 Application to Prostate Cancer Data
Figure 4: Maximum Multiple R vs. model size using the screened genes by the competing methods Air-HOLP, Ridge-HOLP, and SIS
@@ 6 Computational Complexity
Air-HOLP shares the same time complexity as Ridge-HOLP. However, Air-HOLP employs an iterative process to update the initial tuning parameter. We investigate how this iterative process impacts the speed of Air-HOLP.
@@
(a) The logarithm of execution time in milliseconds vs. number of features $p$
(b) Execution time of Air-HOLP vs. Ridge-HOLP in seconds
Figure 5: Compression of the execution time between Ridge-OLS, Ridge-HOLP, and Air-HOLP. On the left, we present the trimmed mean of the logarithm of execution time in milliseconds for $10$ simulations ($10\%$ trimming). On the right, we present a boxplot of the execution time in seconds for $10$ simulations
@@ 7 Discussion
Traditional feature screening methods such as SIS measure the marginal relations between features and the response, which can fail when features are correlated. Ridge-HOLP was developed to address this issue and we have now further built on Ridge-HOLP by the development of Air-HOLP, which incorporates adaptively selecting the ridge tuning parameter. This adaptive selection aims to minimize prediction error, thereby effectively balancing the bias-variance trade-off and enhancing feature screening performance.
Through extensive simulation studies, we demonstrated that Air-HOLP consistently outperforms Ridge-HOLP with a fixed tuning parameter across various simulation settings. Air-HOLP achieves higher sure screening probabilities and lower sure screening thresholds. The advantage of Air-HOLP is more pronounced when the number of samples $n$ is closer to the number of features $p$. Additionally, both Air-HOLP and Ridge-HOLP outperformed SIS in correlated settings, making Air-HOLP especially useful for identifying relevant features in high-dimensional correlated data. However, the ridge penalty is more effective when the correlations between the features are at a similar level. Moreover, Air-HOLP has the same time complexity as Ridge-HOLP and also enjoys being computationally efficient. We also found that in most simulations, Air-HOLP converged as quickly as in $3$ to $6$ iterations. Therefore, we set the maximum number of iterations to be $10$ in our code. We further applied Air-HOLP, Ridge-HOLP and SIS to a prostate cancer genetic dataset. The results confirmed the good performance of Air-HOLP.
While Air-HOLP provides significant advantages for feature screening, certain limitations should be noted. Both Air-HOLP and Ridge-HOLP are designed to measure linear relationships between features and the response and may fail to capture non-linear dependencies. Additionally, these methods are optimized for cases where $p>n$. Although they are functional when $n>p$, their computational efficiency is not ideally suited for such scenarios. These limitations suggest potential areas for future improvement. Moreover, the iterative framework of Algorithm 1, while specifically designed for the ridge penalty, may be extended and adapted to other L2-regularized penalties. Future research could explore these extensions as well as generalise Air-HOLP to non-linear models.
@@ Acknowledgments
Samuel Muller was supported by the Australian Research Council Discovery Project Grant (DP230101908).
Supplementary information
The full simulation results and the code implementing Air-HOLP are made available on GitHub at https://github.com/Logic314/Air-HOLP.git.
Author contributions
Ibrahim Joudah conceived the original research idea, developed the R codes, and conducted the simulation study. Samuel Muller and Houying Zhu provided supervision and guidance throughout the project. All authors collaboratively contributed to developing the research methodology, including refining the Air-HOLP method, designing the simulation study, and applying to real-world data. All authors contributed to writing and revising the manuscript.